OpenStudy (ashchu117):

MEDALS!!!!!!!!!! @pooja195

OpenStudy (anonymous):

;]

pooja195 (pooja195):

post a question -_-

OpenStudy (anonymous):

so whats the question?

OpenStudy (jordan123321):

idk

OpenStudy (anonymous):

@Ashchu117 ! what is the question!?

OpenStudy (ashchu117):

It wouldn't let me click the button sorry :( @pooja195

OpenStudy (anonymous):

ok

pooja195 (pooja195):

its ok :) i dont know this though ;-;

OpenStudy (ashchu117):

but ur a Mathlete XD

OpenStudy (anonymous):

me either

OpenStudy (ashchu117):

@jordan123321 The button wasn't working sorry

OpenStudy (ashchu117):

brb

OpenStudy (ashchu117):

back

OpenStudy (ashchu117):

@e.mccormick @Compassionate @Destinymasha

OpenStudy (ashchu117):

@nincompoop @sammixboo @bibby

OpenStudy (anonymous):

what do you think the answer is?

OpenStudy (ashchu117):

ummm+

OpenStudy (ashchu117):

B

OpenStudy (ashchu117):

hello?

OpenStudy (ashchu117):

:(

OpenStudy (ashchu117):

@arabpride @EclipsedStar @sleepyjess

OpenStudy (anonymous):

Why do you think its B?

OpenStudy (sleepyjess):

eewww completing the square >.< give me a minute to review this, unless @Andrewthehelper knows

OpenStudy (ashchu117):

because they all look like products of the equation

OpenStudy (ashchu117):

@abbychandler I need help

OpenStudy (anonymous):

:P

OpenStudy (sleepyjess):

lol

OpenStudy (ashchu117):

omg XD

OpenStudy (ashchu117):

ok hahaha gimme a sec

OpenStudy (ashchu117):

or 840

OpenStudy (ashchu117):

(the vid is 14mins long and there are 840 seconds in that amount of time)

OpenStudy (ashchu117):

yeah I got lost 2 mins into it @Andrewthehelper @sleepyjess

OpenStudy (lilshane):

if this is not a question plz cloes this

pooja195 (pooja195):

this is a question -_- a MATH one

OpenStudy (ashchu117):

I attached the question as an image @lilshane

OpenStudy (ashchu117):

thx @pooja195

OpenStudy (anonymous):

efforless method.. check from the options about the sum and the product of root technical method... 1. Make the coeff of x^2 = 1 (take out 3 from the whole equation) 2 add and subtract the half of the coeff of x (+9 - 9 ) 3. complete the square you got all the terms

OpenStudy (lilshane):

ok im so sorry @Ashchu117

OpenStudy (ashchu117):

np @lilshane

OpenStudy (ashchu117):

wowowow wait @divu.mkr can you like draw it out or something please cuz thats too much math talk 4 me to understand

OpenStudy (ashchu117):

.-.

OpenStudy (ashchu117):

I'm like a 2nd grade idiot here lol (im really in 8th grade tho and im in Algebra 1 honors)

OpenStudy (anonymous):

give it a shot... i will try to assist you.!

OpenStudy (ashchu117):

ok so if we make the coefficient x^2 then the new equation would look like 3(x^2)2 = 6; x = −3 + square root 2, x = −3 − square root 2.................. right?

OpenStudy (ashchu117):

oops wait

OpenStudy (ashchu117):

3(x^2) = 6; x = -3 + square root 2, x = -3 - square root 2

OpenStudy (ashchu117):

yes or no?

OpenStudy (anonymous):

you checked for sum and product?

OpenStudy (ashchu117):

I don't know how to do that or what that means

OpenStudy (ashchu117):

like I said im an idiot :P

OpenStudy (anonymous):

i am trying to say that the product of the two roots in a quadratic equation is -(coeff of x)/ (coeff of x^2) and product = constant term/(coeff of x^2) give a check for product and sum from the options

OpenStudy (ashchu117):

oops wait i copied the wrong answer XD

OpenStudy (ashchu117):

hold up

OpenStudy (ashchu117):

*wrong equation not answer sorry

OpenStudy (ashchu117):

x^2 + 18x + 15 = 0 <- so there's step one

OpenStudy (ashchu117):

("<-" is an arrow, not part of the equation)

OpenStudy (anonymous):

you dint took 3 out of the whole equation... you just took out from x^2

OpenStudy (ashchu117):

ooohhhh so its x^2 + 15x + 12 = 0

OpenStudy (ashchu117):

or just x^2 + 15x + 15 = 0 ?

OpenStudy (ashchu117):

ok so submitted it and I chose A....... A was wrong................

OpenStudy (ashchu117):

*so I submitted